# Circuit-tuning: A Mechanistic Approach for Identifying Parameter Redundancy and Fine-tuning Neural Networks

### v1

@@ D.1.2 Reasoning-free Tasks
$\mathcal{L}=\mathcal{L}_{NLL}+\beta\cdot|logit(W_{pron}|W_{END})-logit(W_{anti-pron}|W_{END})|$
@@
For evaluation, we use WinoBias (Zhao et al., 2018) which is a classic dataset for coreference resolution focused on gender bias. There are two types of sentences in WinoBias. The Type 1 sentences require world knowledge related to the context to perform coreference resolution, e.g. “The farmer knows [the editor] because [he] is really famous”. The Type 2 sentences can be resolved using syntactic information, e.g. “The CEO called [the hairdresser] and paid [her] over the phone”. In practice, we only use the Type 2 sentences to avoid ambiguity. The test set contains 40 types of occupations in total.
To better evaluate the performance of gender de-biasing, we adopt the concept of prejudice risk from (Liu et al., 2024) which is used to measure the stereotype in large language models. Specifically, given an occupation $x\in X$, a binary gender attribute $y\in\{male,female\}$ and a context $c\in C$, the stereotype of a model $M$ against $x$ about $y$ in the context $c$ is
$s_{y|x}^{M}(c)=\frac{p_{y|x}^{M}(c)}{p_{y|x}^{*}(c)}-1$
@@
$R^{p}=\mathbb{E}_{x\sim X}(r_{x}^{p})$
where $r_{x}^{p}=J(\mathbb{E}_{c\sim C}(s_{y|x}^{M}(c))$ is the prejudice risk of one occupation $x$, and $J(s_{y|x}^{M}(c))=\max\limits_{y\in Y}\{max\{s_{y|x}^{M}(c),0\}\}$ is the discrimination risk criterion. For details, please refer to (Liu et al., 2024).
Reading Comprehension We use SQuAD 2.0 (Rajpurkar et al., 2018) as the dataset. The input contains a paragraph from a passage, and a question related to that paragraph. The answer could be a word or a phrase in the paragraph, or “<No Answer >” which means the answer cannot be found in the paragraph. An example for this task is shown in Table 7.
@@
Paragraph The Normans (Norman: Nourmands; French: Normands; Latin: Normanni) were the people who in the 10th and 11th centuries gave their name to Normandy, a region in France. They were descended from Norse (”Norman” comes from ”Norseman”) raiders and pirates from Denmark, Iceland and Norway who, under their leader Rollo, agreed to swear fealty to King Charles III of West Francia. Through generations of assimilation and mixing with the native Frankish and Roman-Gaulish populations, their descendants would gradually merge with the Carolingian-based cultures of West Francia. The distinct cultural and ethnic identity of the Normans emerged initially in the first half of the 10th century, and it continued to evolve over the succeeding centuries.
@@
D.2 Details for Implementation
We use Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (Dubey et al., 2024) for this task. We set the output of the attention and the MLP in each layer as upstream nodes, and the input of the query, key, value, and the MLP in each layer as downstream nodes. Different from GPT2-small, an MLP layer in Llama is too big to be a node, so we split the input and output of each MLP layer into 64-dimensional MLP heads. Details are shown in Table 8.
Table 8: The settings of the nodes and their corresponding parameters in the complex tasks. For model sizes 1B/3B/8B, the number of layers $L=16/28/32$, and the number of attention heads in each layer $H=32/24/32$, and the number of MLP heads in each layer $H^{*}=128/128/224$. The notations for parameters are specified in Appendix B.
@@ D.2 Details for Implementation
For circuit-tuning, we set the number of edges $N$ to 2000, 3000, and 4000 for 1B/3B/8B models respectively. Since the tasks are much more complex, the value of $N$ is decided by experience. The batch size is set to 16 in all experiments. We set the learning rate to 3e-5 for the mathematics and reading comprehension tasks, and 1e-4 for other tasks. Mini-batch SGD with a momentum equal to 0.9 is used as the optimizer. During training, we perform circuit discovery every 8 steps after optimization for efficiency, which is different from the experiments on GPT2-small in which we perform circuit discovery rightly after an iteration step. For each task, we train the model until performance cannot be further improved.
For LoRA, we set $r=32,\alpha=64$ for all experiments. For full fine-tuning and LoRA, we use the same optimizer as that in circuit-tuning. For all other hyper-parameters such as learning rate, batch size, training steps, and so on, we just sweep over a range of choices and choose the best ones.
In practice, we find that circuit-tuning is much more stable than full fine-tuning and LoRA. When we sweep over a range of hyper-parameters, we notice that full fine-tuning and LoRA are quite sensitive to the change of learning rate, batch size, training steps, and so on. When it comes to circuit-tuning, the change in evaluation result is relatively moderate while still maintaining good performance. Thus it echoes our discussion of the stability of circuit-tuning in Section 4.3.
D.3 Details for Evaluations on General Capabilities
@@
For general capabilities, we use MMLU (Hendrycks et al., 2020), Winogrande (Sakaguchi et al., 2021) and IFEval (Zhou et al., 2023). For MMLU, the evaluation metric is the average accuracy over all categories. For Winogrande, we use the development set for convenience, and the evaluation metric is accuracy. For IFEval which is to test the instruction following ability of a model, each prompt contains one or multiple verifiable instructions, thus the evaluation metric is divided into the prompt-level accuracy and instruction-level accuracy. Due to the randomness of generation, each response is tested under multiple transformations, thus the metric is further divided into strict criterion and loose criterion. In practice, we use the prompt-level and instruction-level accuracy averaged on the strict and loose criteria.
For other capabilities, we consider reasoning, coding, and multilingual capabilities. For reasoning, we use GPQA (Rein et al., 2023) with accuracy as the metric. For coding, we use HumanEval (Chen et al., 2021) with pass@1 as the metric. For multilingual capability, we use MGSM (Shi et al., 2022) with the accuracy averaged on all languages.
For each evaluation benchmark, we test for 5 times and get the averaged result.
D.4 Influence of the Regularization Term on Gender De-biasing
We visualize the prejudice risk of the models before and after gender de-biasing in Figure 12. Note that the coefficient for regularization $\beta=0.5$. Since there are only 40 types of occupations in the test set, we regard the distribution of prejudice risk as normal distribution and perform interpolation on the computed results. For convenience, we only show the results of GPT2-small and Llama-3.2-1B-instruct in the figure.
The dynamic process of de-biasing can be observed from right to left. It is obvious that with a regularization term in the loss function, the distribution of the prejudice risk in occupations is more concentrated to a smaller value. The results demonstrate the effectiveness of our method in modifying a model’s stereotype, providing new insights in aligning AI systems with human values.
@@ D.4 Influence of the Regularization Term on Gender De-biasing
Figure 12: The comparison between biased and de-biased models.

### v2

@@ G.1.2 Reasoning-free Tasks
$\mathcal{L}=\mathcal{L}_{NLL}+\beta\cdot|logit(W_{pron}|W_{END})-logit(W_{anti-pron}|W_{END})|$ (32)
@@
For evaluation, we use WinoBias [65] which is a classic dataset for coreference resolution focused on gender bias. There are two types of sentences in WinoBias. The Type 1 sentences require world knowledge related to the context to perform coreference resolution, e.g. “The farmer knows [the editor] because [he] is really famous". The Type 2 sentences can be resolved using syntactic information, e.g. “The CEO called [the hairdresser] and paid [her] over the phone". In practice, we only use the Type 2 sentences to avoid ambiguity. The test set contains 40 types of occupations in total.
To better evaluate the performance of gender de-biasing, we adopt the concept of prejudice risk from [33] which is used to measure the stereotype in large language models. Specifically, given an occupation $x\in X$, a binary gender attribute $y\in\{male,female\}$ and a context $c\in C$, the stereotype of a model $M$ against $x$ about $y$ in the context $c$ is
$s_{y|x}^{M}(c)=\frac{p_{y|x}^{M}(c)}{p_{y|x}^{*}(c)}-1$ (33)
@@
$R^{p}=\mathbb{E}_{x\sim X}(r_{x}^{p})$ (34)
where $r_{x}^{p}=J(\mathbb{E}_{c\sim C}(s_{y|x}^{M}(c))$ is the prejudice risk of one occupation $x$, and $J(s_{y|x}^{M}(c))=\max\limits_{y\in Y}\{max\{s_{y|x}^{M}(c),0\}\}$ is the discrimination risk criterion. For details, please refer to [33].
Reading Comprehension We use SQuAD 2.0 [48] as the dataset. The input contains a paragraph from a passage, and a question related to that paragraph. The answer could be a word or a phrase in the paragraph, or “<No Answer >" which means the answer cannot be found in the paragraph. An example for this task is shown in Table 7.
@@
Paragraph The Normans (Norman: Nourmands; French: Normands; Latin: Normanni) were the people who in the 10th and 11th centuries gave their name to Normandy, a region in France. They were descended from Norse ("Norman" comes from "Norseman") raiders and pirates from Denmark, Iceland and Norway who, under their leader Rollo, agreed to swear fealty to King Charles III of West Francia. Through generations of assimilation and mixing with the native Frankish and Roman-Gaulish populations, their descendants would gradually merge with the Carolingian-based cultures of West Francia. The distinct cultural and ethnic identity of the Normans emerged initially in the first half of the 10th century, and it continued to evolve over the succeeding centuries.
@@
G.2 Details for Implementation
We use Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct [11] for this task. We set the output of the attention and the MLP in each layer as upstream nodes, and the input of the query, key, value, and the MLP in each layer as downstream nodes. Different from GPT2-small, an MLP layer in Llama is too big to be a node, so we split the input and output of each MLP layer into 64-dimensional MLP heads. Details are shown in Table 8.
Table 8: The settings of the nodes and their corresponding parameters in the complex tasks. For model sizes 1B/3B/8B, the number of layers $L=16/28/32$, and the number of attention heads in each layer $H=32/24/32$, and the number of MLP heads in each layer $H^{*}=128/128/224$. The notations for parameters are detailed in Appendix C.
@@ G.2 Details for Implementation
For circuit-tuning, we set the number of edges $N$ to 2000, 3000, and 4000 for 1B/3B/8B models respectively. The choice of $N$ is considered with the size of the models and the complexity of the tasks. In practice, we tried several values of $N$ and found that the task performance generally converges at these values. Actually, the value of $N$ corresponds to the intrinsic dimension which is the minimal dimensionality required in a solution space. If the chosen value is smaller than that, then the result may not be optimal. If the value of exceeds that, then the improvement in the final result may be quite small, which is demonstrated in Section 5. In fact, the choice of $N$ is quite similar to the choice of rank $r$ in LoRA, except that our method is able to localize the key parameters more accurately.
The batch size is set to 16 in all experiments. We set the learning rate to 3e-5 for the mathematics and reading comprehension tasks, and 1e-4 for other tasks. Mini-batch SGD with a momentum equal to 0.9 is used as the optimizer. During training, we perform circuit discovery every 8 steps after optimization for efficiency, which is different from the experiments on GPT2-small in which we perform circuit discovery rightly after an iteration step. For each task, we train the model until performance cannot be further improved.
For LoRA, we set $r=32,\alpha=64$ for all experiments, since this is the best setting we could get. We have swept a wide range of values for rank $r$ and alpha $\alpha$, and find that the performance cannot be further improved or even decreases when $r$ increased over 32. Actually, Hu et al. [20] found similar phenomenon when they studied LoRA. From our opinion, this is because LoRA falis to accurately figure out the key parameters needed to be fine-tuned, since all the parameters are changed after LoRA fine-tuning.
For full fine-tuning and LoRA, we use the same optimizer as that in circuit-tuning. For all other hyper-parameters such as learning rate, batch size, training steps, and so on, we just sweep over a range of choices and choose the best ones.
In practice, we find that circuit-tuning is much more stable than full fine-tuning and LoRA. When we sweep over a range of hyper-parameters, we notice that full fine-tuning and LoRA are quite sensitive to the change of learning rate, batch size, training steps, and so on. When it comes to circuit-tuning, the change in evaluation result is relatively moderate while still maintaining good performance.
Experiments are conducted on 8 $\times$ A800 Nvidia GPUs.
G.3 Details for Evaluations on General Capabilities
@@
For general capabilities, we use MMLU [19], Winogrande [52] and IFEval [66]. For MMLU, the evaluation metric is the average accuracy over all categories. For Winogrande, we use the development set for convenience, and the evaluation metric is accuracy. For IFEval which is to test the instruction following ability of a model, each prompt contains one or multiple verifiable instructions, thus the evaluation metric is divided into the prompt-level accuracy and instruction-level accuracy. Due to the randomness of generation, each response is tested under multiple transformations, thus the metric is further divided into strict criterion and loose criterion. In practice, we use the prompt-level and instruction-level accuracy averaged on the strict and loose criteria.
For other capabilities, we consider reasoning, coding, and multilingual capabilities. For reasoning, we use GPQA [49] with accuracy as the metric. For coding, we use HumanEval [7] with pass@1 as the metric. For multilingual capability, we use MGSM [55] with the accuracy averaged on all languages.
To check if the general capabilities as well as other capabilities are affected after fine-tuning, we compute the relative change in performance for each capability. For example, if the evaluation results before and after fine-tuning are $x$ and $x^{\prime}$, then the relative change is $\frac{x^{\prime}-x}{x}$. In practice, we test each model on each benchmark for 10 times. We find that the relative performance change is varies little among different settings, which further shows the robustness of circuit-tuning. Therefore the final relative change in performance of each capability is averaged over different sizes of models and fine-tuning tasks, as discussed in Section 5.2, and the standard errors are also the averaged results. The final results are shown in Figure 6. It can be seen that circuit-tuning preserves general capabilities better than others, demonstrating its ability to locate the key components precisely for a specific task.
G.4 Influence of the Regularization Term on Gender De-biasing
We selectively add $\beta\cdot|\mathcal{L}_{m}|$ as a regularization term in the loss $\mathcal{L}$ for gender de-biasing to improve performance, as discussed in Appendix G.1.2. We visualize the prejudice risk of the models before and after gender de-biasing in Figure 6. Note that the coefficient for regularization $\beta=0.5$. Since there are only 40 types of occupations in the test set, we regard the distribution of prejudice risk as normal distribution and perform interpolation on the computed results. For convenience, we only show the results of GPT2-small and Llama-3.2-1B-instruct in the figure.
The dynamic process of de-biasing can be observed from right to left in Figure 6. It is obvious that with a regularization term in the loss function, the distribution of the prejudice risk in occupations is more concentrated to a smaller value. The results demonstrate the effectiveness of our method in modifying a model’s stereotype, providing new insights in aligning AI systems with human values.